OpenStudy (anonymous):

What is 42% of 150? Will you guys show me how to do it? Help!

OpenStudy (tkhunny):

150 is 100% of 150 -- Prove it! 150*1.00 = 150 75 is 50% of 150 -- Prove it! 150*0.50 = 75 37½ is 25% of 150 -- Prove it! 150*0.25 = 37.5 It would behoove all others to read the Code of Conduct and heed its wisdom.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (tkhunny):

Come on. You can do it with the hints I provided. Read it carefully and reason it out.

OpenStudy (amistre64):

1% of n is simply, n/100 once you know 1% you can find any %

OpenStudy (anonymous):

How would you set up the problem..?

OpenStudy (anonymous):

I don't get how to set up the problem

OpenStudy (tkhunny):

I gave you three examples. Do that.

OpenStudy (anonymous):

Like: 150*0.42?

OpenStudy (anonymous):

42% = .42 .42 * 150 = you answer

OpenStudy (anonymous):

63?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

your welcome